# radarODE: An ODE-Embedded Deep Learning Model for Contactless ECG Reconstruction from Millimeter-Wave Radar

### v1

@@ III-B2 Signal Pre-Processing with Synchrosqueezed Wavelet Transform (SST)
Based on the proposed radar signal model in (12), the next step is to localize when the vibrations happen (i.e., $T_{1}$, $T_{2}$). Figure 1 shows that the high SNR radar signal could reveal prominent peaks of $v_{1}$ and $v_{2}$ in the time domain, but in most cases, these two peaks (especially $v_{2}$) could be ruined by the noise. Therefore, this research decides to extract the time-frequency domain information from the spectrogram obtained by synchrosqueezed wavelet transform (SST) [44], and the two vibrations can then be localized by the SCEG module proposed later in Section III-C2.
@@
where $\psi^{*}$ is the complex conjugate of the chosen mother wavelet, and $a$, $b$ are the adjustable scaling and translation factors for the wavelet $\psi$ to extract frequency- and time-domain information, respectively. In this research, the Morlet wavelet is selected as mother wavelet because it is widely used for vibration signal processing, especially for time-frequency localization [45, 46].
@@
where $A(b)=\left\{a;W_{s}(a,b)\neq 0\right\}$, and $\delta$ represents the Dirac-delta function in a distribution version to smoothly squeeze the spread-out energy into a narrow band around the instantaneous frequency [44].
The resultant spectrograms obtained from the same piece of radar signal (Figure 4(a)) using different time-frequency representation tools are shown in Figure 4. The spectrogram yielded by STFT in Figure 4(b) could show the rough positions of each $v_{1}$ but cannot clearly reveal the position for $v_{2}$. In comparison, CWT gives a sharp position for both vibrations, but the energy is still spread out and sometimes mixed with the background, as shown in Figure 4(c). By further concentrating the energy distribution, the spectrogram obtained by SST has a relatively clean background and sharp peaks for the vibrations as shown in Figure 4(d), reducing the burden of the deep-learning-based SCEG module in extracting latent features (i.e., $T_{1}$, $T_{2}$).
Figure 4: Spectrograms obtained from the same radar signal: (a) Radar signal $\tilde{x}(t)$; (b) STFT result; (c) CWT result; (d) SST result with $T_{1}$, $T_{2}$ labeled.
@@
III-C1 Single Cycle ECG Length Estimation
The first module of radarODE aims to estimate the length of each single cycle ECG piece by calculating the interval between two consecutive heartbeats (i.e., PPI), and these heartbeats will bring intensive energy into the received radar signal [21]. Therefore, the energy plots $\hat{x}(t)$ of the radar signal $\tilde{x}(t)$ can be obtained by simply adding the SST spectrogram along the frequency axis and are used for peak detection as shown in Figure 5a. In addition, it is found in practice that the algorithms designed for ECG R-peaks detection are also applicable for the peaks detection from the radar energy plot $\hat{x}$ (omit $(t)$ for simplicity) [47], because apparently the energy plot shares a similar upsurge with the ECG signal. However, the $50$ synchronous radar signals provided in the dataset ($\{\tilde{x}_{1},\ldots,\tilde{x}_{50}$}) do not always have high SNR, potentially causing the wrong detection as shown in Figure 5b. Therefore, a new algorithm is proposed as in Algorithm 1 to eliminate the wrong detection obtained from $50$ series of synchronous radar energy plots ${X}_{\mathcal{L}}$, with the length of each $\hat{x}_{i}\in{X}_{\mathcal{L}}$ equaling to $l$.
The design of Algorithm 1 is based on the fact that the PPI for healthy people tends to be unchanged in adjacent cardiac cycles. In this case, the long-term radar energy plots are firstly sliced into short segments as shown in the Initialization stage in Algorithm 1, and then the biopeaks algorithm [48] implemented in NeuroKit2 [47] is used for detecting all the potential peaks $P$ from each energy plot segment $\hat{x}^{j}_{i}$ as:
@@ III-C1 Single Cycle ECG Length Estimation
$\hat{f}(p)=\frac{1}{nh}\sum_{c=1}^{n}K\left(\frac{p-PPI_{c\in\mathcal{C}}}{h}\right)$ (18)
where $\hat{f}(p)$ means the estimated probability density function at point $p$, $n$ is the number of all the estimated PPI in $PPI_{\mathcal{C}}$, $K$ represents the Gaussian kernel function and $h=n^{-1/5}$ is the bandwidth of the kernel. Lastly, the final PPI for the current segment is selected as argument $p$ when $\hat{f}(p)$ achieves maximum as shown in Line 15 in Algorithm 1, and the long-term PPI estimation can be obtained step by step as the Algorithm 1 terminated.
@@
Remark 2. Algorithm 1 actually provides an accurate PPI estimation for coarse cardiac monitoring tasks as demonstrated later in Section IV-C1, and can be directly used for the monitoring of heart rate and HRV.
III-C2 Single Cycle ECG Generator (SCEG)
Based on the yielded $PPI$, the SST spectrogram can be sliced into segments corresponding to a single cardiac cycle, and the aim of the SCEG module is to reconstruct the ECG for each single cardiac cycle, hence realizing the transformation from mechanical to electrical domain. In general, the input of the SCEG is $N$ segments of the SST plot within $[1,25]$ Hz with the size of $F\times T$ in frequency and time axis, and the output is the corresponding $N$ pieces of ECG with the same length $T$, as shown in Figure 6. In practice, the deep neural network only accepts the inputs/outputs with the same size. Therefore, the actual SST segment is centered at the current cardiac cycle and expands to $4$ seconds, and the corresponding ECG ground truth is resampled to a fixed length of $200$ for loss calculation.
For the architecture design, the SCEG module adopts the popular backbone, encoder, and decoder structure as verified by enormous image-related tasks [50, 51, 52], with detailed parameters shown in Table I. In addition, a feature fusion block is added after the decoder to fuse the temporal and morphological features and generate the final ECG reconstructions. The detailed implementation of each part in Table Iwith explanations can be elaborated as:
Figure 6: Architecture of SCEG with SST segments as input and single cycle ECG pieces as output.
TABLE I: Parameters used for SCEG
@@ III-C2 Single Cycle ECG Generator (SCEG)
Output single cycle ECG piece $(N,1,200)$
@@ 1. $C_{in}$: Input channel, $C_{out}$: Output channel, $K$: Kernel size, $S$: Stride
    According to the literature, the ResNet network with residual blocks is widely used as the backbone for feature extraction in many fields [53, 54, 55], and this research will use a similar structure with $4$ layers of residual blocks and downsample blocks as shown in Figure 7, with the key parameters listed in Table I. In addition, the traditional 2D convolution is all replaced by deformable 2D convolution [56] with deformable kernels (instead of square kernels) to fit the irregular shape of the target patterns, as shown in red triangles in Figure 6.
@@
    Squeezer and Encoder: The output feature map from the backbone should involve both frequency- and time-domain features, and the squeezer is simply used to squeeze the frequency-domain feature using 2D convolution (Conv2d) and output the feature map with the temporal feature only. Then, the encoder assembled by three 1D transposed convolution (Transconv1d) blocks is used to further extract the temporal feature, with each block comprising Transconv1d, BN and ReLU. Finally, the output feature map from the encoder should contain the latent information about the $v_{1}$, $v_{2}$ (i.e., ($T_{1}$, $T_{2}$)).
@@
    In this case, the ODE decoder is designed as a branch to assist the transformation between cardiac mechanical and electrical activities (radar to ECG) by introducing two prior facts in the real world:
    • Although the biological one-to-one correspondence between cardiac electrical and mechanical activities (excitation–contraction coupling) is too complex to be modeled, it is safe to claim that the measurements of mechanical activities generally lag behind those of electrical activities with a short time delay $\tau$ [31, 32].
    • The shape of the ECG piece can be modeled morphologically using ODEs without any biological/chemical knowledge [58], and the introduction of ODEs could provide morphological prior to accelerating the convergence of the model training process and also providing extra robustness in the presence of noises.
    Inspired by the above two facts, the ODE decoder is designed as a parameters estimation part and an ODE solver as shown in Figure 6, and the solution of the ODEs will be shifted to the left with time $\tau$. To be specific, the parameters estimation part contains four linear blocks (Linear Layer, BN, Tanh) to project the latent space yielded by the initial decoder into parameters $\eta$, $\tau$, and $\eta$ will be sent to an ODE solver to solve a 3D trajectory denoted by $(x,\ y,\ z)$ as
@@
    where $\mathcal{F}$ represents five characteristic peaks (PQRST) in a single cycle ECG signal, and the whole ODEs can be interpreted as manipulating each peak along a unit circle by varying the value of $\eta=\{a_{e_{f}},b_{e_{f}},\theta_{e_{f}}\}$ to adjust corresponding amplitude, width and position of each peak. After specifying all $15$ parameters $\eta$ ($3$ for each peak) and the initial conditions of $(x,y,z)$, the value of $z$ can be solved out by the ODE solver using Euler method to get the final single cycle ECG signal as the morphological feature. In practice, the default values for $\eta$ are provided in advance as in Table II, and the estimated parameters within the range of $[-1,1]$ are used to scale the default values.
    TABLE II: Default values for $\eta$
    $a_{e_{f}}$ $5$ $-100$ $480$ $-120$ $8$
    $b_{e_{f}}$ $0.25$ $0.1$ $0.1$ $0.1$ $0.4$
@@
    Feature Fusion: The feature fusion module could leverage respective advantages of the morphological and temporal features and generate the final ECG signal for loss calculation, because the morphological feature only focuses on five peaks to provide a rough shape of ECG with calibrated peaks, and the temporal feature reserves all the other feature neglected in (19) to help the final reconstruction [59]. Therefore, two features are first fused together by multiplication into one and then stacked four times by itself. Then, the stacked feature is encoded and decoded as in Table I to yield the final single cycle ECG piece.
The last step of SCEG is to resample all the ECG pieces generated after the feature fusion part with respect to the previous PPI estimation, and the resampled ECG pieces are concatenated in time sequence to form the final morphological reference for long-term ECG reconstruction.
@@ III-C3 Long-Term ECG Reconstruction
The long-term ECG reconstruction network adopts a simpler encoder-decoder-fusion structure than SCEG as shown in Figure 2, because the morphological reference reduces the burden of learning the ECG peaks and patterns. In addition, the input changes to arbitrary segments of $50$ synchronous time-domain radar signal instead of the SST segments within a single cardiac cycle. Therefore, the encoder for raw radar signal has the same structure as the backbone in SCEG, only with the deformable 2D convolution replaced by 1D convolution with $K=5$, $S=2$. Then, the decoder is performed by four Transconv1d blocks with the same $K$ and $S$ as an encoder to generate the temporal feature, and the feature fusion module also shares the same strategy as SCEG. Lastly, $9$ stacks of the temporal convolutional network (TCN) with dilation fact of $2$ are added as suggested in [35] to recursively generate the ECG based on the historical reconstructions.
IV Experimental and Result Evaluation
IV-A Dataset and Benchmark Introduction

### v2

@@ III-B2 Signal Pre-Processing with Synchrosqueezed Wavelet Transform (SST)
Based on the proposed radar signal model in (12), the next step is to enhance the prominent vibrations (i.e., $T_{1}$, $T_{2}$). Figure 1 shows that the high SNR radar signal could reveal prominent peaks of $v_{1}$ and $v_{2}$ in the time domain, but in most cases, these two peaks (especially $v_{2}$) could be ruined by noise. Therefore, this research decides to extract the time-frequency domain information from the spectrogram obtained by synchrosqueezed wavelet transform (SST) [37], and the two vibrations can then be localized by the SCEG module proposed later in Section III-C2.
@@
where $\psi^{*}$ is the complex conjugate of the chosen mother wavelet, and $a$, $b$ are the adjustable scaling and translation factors for the wavelet $\psi$ to extract frequency- and time-domain information, respectively. In this research, the Morlet wavelet is selected as mother wavelet because it is widely used for vibration signal processing, especially for time-frequency localization [38].
@@
where $A(b)=\left\{a;W_{\tilde{x}}(a,b)\neq 0\right\}$, and $\delta$ represents the Dirac-delta function in a distribution version to smoothly squeeze the spread-out energy into a narrow band around the instantaneous frequency [37].
The quality of the resultant spectrograms can be evaluated using power spectrogram entropy (PSE) [33], with a small value indicating that the energy is concentrated around a certain frequency. The calculated PSE for the spectrogram produced by STFT, CWT and SST is $0.94$, $0.90$ and $0.76$ respectively, with the visualized results shown in Figure LABEL:sub@fig:compare_stftLABEL:sub@fig:compare_cwt and LABEL:sub@fig:compare_sst. It is clear that the spectrogram obtained by STFT only shows the rough positions of each $v_{1}$, while CWT gives a sharp position for both vibrations, but the energy is still spread out. By further concentrating the energy distribution, SST produces the spectrogram with a relatively clean background and sharp peaks for the vibrations, reducing the burden of the deep-learning-based SCEG module in extracting latent features (i.e., $T_{1}$, $T_{2}$).
Figure 3: Spectrograms obtained from the same radar signal: (a) Radar signal $\tilde{x}(t)$; (b) STFT result; (c) CWT result; (d) SST result with $T_{1}$, $T_{2}$ labeled, revealing concentrated energy around vibrations and clean background.
@@
III-C1 Single-cycle ECG Length Estimation
The first module of radarODE aims to estimate the length of each single-cycle ECG piece by calculating the interval between two consecutive heartbeats (i.e., PPI) from the energy plots $\hat{x}$ of the radar signal $\tilde{x}$ (omit $(t)$ for simplicity) as shown in Figure LABEL:sub@fig:correct_detection. The energy plot is obtained by simply adding the spectrogram along frequency axis, but the peak detection results may not be promising due to low SNR signals as shown in Figure LABEL:sub@fig:wrong_detection. Therefore, a new algorithm is proposed as in Algorithm 1 to eliminate the wrong detection obtained from $50$ radar energy plots ${X}_{\mathcal{L}}$, with the length of each $\hat{x}_{i}\in{X}_{\mathcal{L}}$ equal to $l$.
The design of Algorithm 1 is based on the fact that the PPI for healthy people tends to be unchanged in adjacent cardiac cycles. In this case, the long-term radar energy plots are firstly sliced into short segments as shown in the Initialization stage in Algorithm 1, and then the biopeaks algorithm implemented in NeuroKit2 [39] is used for detecting all the potential peaks $P$ from each energy plot segment $\hat{x}^{j}_{i}$ as:
@@ III-C1 Single-cycle ECG Length Estimation
$KDE:\ \hat{f}(p)=\frac{1}{nh}\sum_{c=1}^{n}K\left(\frac{p-PPI_{c\in\mathcal{C}}}{h}\right)$ (18)
where $\hat{f}(p)$ means the estimated probability density function at point $p$, $n$ is the number of all the estimated PPI in $PPI_{\mathcal{C}}$, $K$ is the Gaussian kernel function and $h=n^{-1/5}$ is the bandwidth of the kernel. Lastly, the final PPI for the current segment is selected as the argument $p$ when $\hat{f}(p)$ achieves the maximum as in Line 15 in Algorithm 1, and the long-term PPI estimation can be obtained step by step as Algorithm 1 terminated.
@@
III-C2 Single-cycle ECG Generator (SCEG)
Based on the yielded $PPI$, the SST spectrogram can be sliced into segments corresponding to a single cardiac cycle, and the aim of the SCEG module is to reconstruct the ECG for each single cardiac cycle, hence realizing the transformation from mechanical to electrical domain. In general, the input of the SCEG is $N$ segments of the SST plot within $[1,25]$ Hz with the size of $F\times T$ on frequency and time axis, and the output is the corresponding $N$ ECG pieces with the same length $T$, as shown in Figure 5. In practice, the deep neural network only accepts the inputs/outputs of the same size. Therefore, the actual SST segment is centered at the current cardiac cycle and expands to $4$ seconds, and the corresponding ECG ground truth is resampled to a fixed length of $200$ for loss calculation.
For architecture design, the SCEG module adopts the popular backbone-encoder-decoder structure as verified by enormous image-related tasks [41, 42], with detailed parameters shown in Table I. In addition, a feature fusion block is added after the decoder to fuse the temporal and morphological features and generate the final ECG reconstructions. The detailed implementation of each part in Table I with explanations can be elaborated as:
Figure 5: Architecture of SCEG with SST segments as input and single-cycle ECG pieces as output.
TABLE I: Structure and Parameters for SCEG
@@ III-C2 Single-cycle ECG Generator (SCEG)
Output single-cycle ECG piece $(N,1,200)$
@@ 1. $C_{in}$: Input channel, $C_{out}$: Output channel, $K$: Kernel size, $S$: Stride
    According to the literature, ResNet with residual blocks is widely used as the backbone for feature extraction in many fields [43], and this research will use a similar structure with $4$ layers of residual blocks and downsample blocks as shown in Figure 6, with the key parameters listed in Table I. In addition, the traditional 2D convolution is all replaced by deformable 2D convolution [44] with deformable kernels (instead of square kernels) to fit the irregular shape of the target patterns, as shown in red triangles in Figure 5.
@@
    Squeezer and Encoder: The output feature map from the backbone should involve both frequency- and time-domain features, and the squeezer is simply used to squeeze the frequency-domain feature using 2D convolution (Conv2d) and output the feature map with the temporal feature only. Then, the encoder assembled by three 1D transposed convolution (Transconv1d) blocks is used to further extract the temporal feature, with each block comprising Transconv1d, BN and ReLU. Finally, the output feature map from the encoder should contain latent information about $v_{1}$, $v_{2}$ (i.e., ($T_{1}$, $T_{2}$)).
@@
    TABLE II: Default values for $\eta$
    $a_{e_{f}}$ $5$ $-100$ $480$ $-120$ $8$
    $b_{e_{f}}$ $0.25$ $0.1$ $0.1$ $0.1$ $0.4$
    In this case, the ODE decoder is designed as a branch to assist the transformation between cardiac mechanical and electrical activities. The main obstacles in modeling such domain transformation are the lack of (a) A compact model for ECG signal and (b) A corresponding explanation between parameters in describing radar signal and ECG signal, e.g., what is the relationship between R peak and $v_{1}$ in Figure 1. In this work, the aforementioned two obstacles can be solved from the following perspectives:
    • The shape of the ECG piece can be modeled morphologically using ODEs in a compact form without any biological/chemical knowledge [46].
    • The measurements of mechanical activities generally lag behind those of electrical activities with a short time delay $\tau$ [26, 27].
    Inspired by the above facts, the ODE decoder is designed as a parameter estimation part and an ODE solver as shown in Figure 5, and the solution of the ODEs will be shifted to the left with time $\tau$. In this manner, the latent information in describing radar signals is first transformed to the parameters for the ECG signal, and then the ODE solver will generate morphological-prior to accelerate the convergence of the model training process and provide extra robustness against noises.
    To be specific, the parameters estimation part contains four linear blocks (Linear Layer, BN, Tanh) to project the latent space yielded by the initial decoder into parameters $\eta$, $\tau$, and $\eta$ will be sent to an ODE solver to solve a 3D trajectory denoted by $(x,\ y,\ z)$ as
@@
    where $\mathcal{F}$ represents five characteristic peaks (PQRST) in a single-cycle ECG signal, and the whole ODEs can be interpreted as manipulating each peak along a unit circle by varying the value of $\eta=\{a_{e_{f}},b_{e_{f}},\theta_{e_{f}}\}$ to adjust corresponding amplitude, width and position of each peak. After specifying all $15$ parameters $\eta$ ($3$ for each peak) and the initial conditions of $(x,y,z)$, the value of $z$ can be solved by the ODE solver using the Euler method to get the final single-cycle ECG signal as the morphological feature. In practice, the default values for $\eta$ are provided in advance as in Table II, and the estimated parameters within the range of $[-1,1]$ are used to scale the default values.
@@
    Feature Fusion: The feature fusion module could leverage respective advantages of the morphological and temporal features and generate the final ECG signal for loss calculation, because the morphological feature only focuses on five peaks to provide a rough shape of ECG with calibrated peaks, and the temporal feature reserves all the other feature neglected in (19) to help the final reconstruction [4]. Therefore, two features are first fused together by multiplication (Mul.) into one and then stacked four times by itself. Then, the stacked feature is encoded and decoded as in Table I to produce the final single-cycle ECG piece.
The last step of SCEG is to resample all the ECG pieces generated after the feature fusion part with respect to the previous PPI estimation, and the resampled ECG pieces are concatenated in a time sequence to form the final morphological reference for long-term ECG reconstruction.
TABLE III: Structure and Parameters for Long-term ECG Reconstruction
Layers Parameters Output Shape
($C_{in}$, $C_{out}$, $K$, $S$)1 $N$: Batch Size
Input raw radar Signal $(N,50,800)$
a. Encoder
    Residual Block $(50,128,5,1)$ $(N,128,800)$
    Downsample Block $(128,128,5,2)$ $(N,128,400)$
    Residual Block $(128,256,5,1)$ $(N,256,400)$
    Downsample Block $(256,256,5,2)$ $(N,256,200)$
    Residual Block $(256,512,5,1)$ $(N,512,200)$
    Downsample Block $(512,512,5,2)$ $(N,512,100)$
b. Decoder
    Transconv1d Block $(512,128,5,2)$ $(N,128,200)$
    Transconv1d Block $(128,16,5,2)$ $(N,16,400)$
    Transconv1d Block $(16,1,5,2)$ $(N,1,800)$
c. Feature Fusion (TCN)
    Feature Stack2 -- $(N,2,800)$
    Dilated Conv1d ×$\times 9$ $K=3$, $D=2$ $(N,1,800)$
Output long-term ECG $(N,1,800)$
2. Stack with the morphological feature as in Figure 2(c).
Figure 7: The structure of the $9$-layer TCN with dilation factor $D=2$ and kernel size $K=3$.
@@ III-C3 Long-Term ECG Reconstruction
The long-term ECG reconstruction network adopts a similar encoder-decoder-fusion structure with SCEG as shown in Figure 2(c), with the detailed structure and parameters shown in Table III and Figure 7. The encoder takes $50$ series of 4-sec time-domain radar signals as input and is composed of three groups of residual and downsample blocks with the deformable 2D convolution in Figure 6 replaced by 1D convolution. Then, the decoder is realized by three Transconv1d blocks to generate the temporal feature.
The temporal feature generated from decoder in Table III will be stacked with morphological feature as illustrated in Figure 2(c), and these two features act as two channels for later feature fusion by the temporal convolutional network (TCN). In general, TCN adopts dilation 1D convolution (Dilated Conv1d) to process multi-channel data structure using expanded receptive field with gaps between elements as shown in Figure 7, and the feature fusion is achieved during channel reduction as a common technique in traditional convolution neural network [38]. Specifically, the solid line in Figure 7 shows the connection of a $4$-layer TCN with the gaps for each layer $l$ as $d=D^{l-1}-1$, and the output $y_{T}$ is predicted based on a receptive field of $15$ input feature points $\{a_{T-14},\dots,a_{T}\}$. In this paper, $9$-layer Dilated Conv1d is adopted with dilation factor $D=2$ and kernel size $K=3$, and the receptive field is $511$ to make the most of contextual information contained in the temporal and morphological features.
IV Dataset and Implementation Details
IV-A Hardware and Environment Settings for Data Collection
The public dataset can be requested from [28] and is collected by the TI AWR-1843 radar with $77$ GHz start frequency and $3.8$ GHz bandwidth, providing good SNR and resolution to detect subtle vibrations that fit the proposed signal model and framework. To realize the 3D beamforming to extract cardiac features from real 3D space, $3$ transmitters (Tx) and $4$ receivers (Rx) are enabled with time division multiplexing multi-input multi-output (TDM-MIMO) applied in chirp transmitting and receiving, and the essential parameters in radar configuration setting are listed in Table IV.
TABLE IV: Parameters for Data Collection Interface
Parameter Value     Parameter Value
Start Frequency $77$ GHz     Frequency Slope $65$ MHz/$\mu$s
Idle Time $10$ $\mu$s     Tx Start Time $1$ $\mu$s
ADC Start Time $6$ $\mu$s     ADC Samples $200$
Sample Rate $5000$ kbps     Ramp End Time $60$ $\mu$s
Rx Gain $30$ dB     Rx Gain Target $30$ dB
Start/End Chirp Tx $0/2$     No. of Chirp Loops $2$
No. of Frames $3600$     Frame Periodicity $50$ ms
The data collection is performed for subjects lying on the bed with quasi-static status to ensure good SNR with the least RBM noise. In addition, radar is placed right above the human chest region in a range of $0.4-0.5$m with minor propagation attenuation as shown in Figure 8, and hence the large- or small-scale signal variations (e.g., path loss, multi-path fading) are not considered in [28].
Figure 8: Environment settings for data collection from quasi-static subject [28].
IV-B Link Budget Analysis
Link budget analysis is a common evaluation for the performance of a radar system by accounting for all gains and losses from the transmitter to the receiver [7, 36], but such analysis is not provided in [28]. For a radar system, the received power $P_{\mathrm{R}}$ can be expressed with respect to the transmitted power $P_{\mathrm{T}}$ as:
$P_{\mathrm{R}}=\frac{G_{\mathrm{T}}G_{\mathrm{R}}\lambda^{2}\sigma}{(4\pi)^{3}R^{4}}P_{\mathrm{T}}$ (21)
where $G_{T}$/$G_{R}$ is the gain of the transmitter/receiver, $\sigma$ is the effective radar cross section (RCS) for the human chest region, and $R$ is the distance between radar and chest [7].
In addition, the lowest detectable signal power of the receiver is defined in [36] as:
$P_{\mathrm{r},\min}=-174\ \mathrm{dBm}+10\log_{10}(B)+\mathrm{NF}+\mathrm{SNR}_{\min}$ (22)
where $-174$ represents the thermal-noise level, NF means the noise figure, $B$ is the bandwidth of the receiver and $\mathrm{SNR}_{\min}$ is the desired SNR considering the afterward signal processing.
By combining (21) and (22), the maximum detectable range can be calculated as:
$R_{\text{max}}=\sqrt[4]{\frac{G_{\mathrm{T}}G_{\mathrm{R}}\lambda^{2}\sigma P_{\mathrm{T}}}{(4\pi)^{3}\ P_{\mathrm{r},\min}}}$ (23)
All the radar-related values in (23) are shown in Table V according to the datasheet of TI AWR-1843 radar [47]. According to the previous work [36, 28], the RCS value can be estimated as $-20$ dBsm for the quasi-static subjects wearing electrically thin cloth (i.e., the thickness is much smaller than the wavelength), with respiration noise filtered during the pre-processing stage. After substituting the values, the lowest detectable signal power is $P_{\mathrm{r},\min}=-53$ dBm, and the maximum detectable range is $R_{\text{max}}=4$ m, revealing that the parameter setting used in [28] could provide the received signal with good SNR.
TABLE V: Parameters for Link Budget Analysis
Parameter Value     Parameter Value
Tx Gain ($G_{T}$) $10$ dBi     Rx Gain ($G_{R}$) $30$ dBi
Tx power ($P_{\mathrm{T}}$) $12$ dBm     Noise Figure (NF) $15$ dB
Wavelength ($\lambda$) $3.9$ mm     Bandwidth (B) $3.8$ GHz
RCS ($\sigma$) $-20$ dBsm     Desired SNR ($\mathrm{SNR}_{\min}$) $10$ dB
IV-C Dataset Description